OpenStudy (anonymous):

[5z+3]+9=12 [5z+3]+9=12 @Mathematics

OpenStudy (anonymous):

5z = 12 - 3 -9 5z = 0 z = 0

OpenStudy (anonymous):

5z+12=12 5z=0 z=0